OpenStudy (anonymous):

Change the given equation into slope-intercept form. 4x - 3y = -12

OpenStudy (texaschic101):

slope intercept form : y = mx + b 4x - 3y = - 12 -3y = -4x - 12 y = (-4x - 12)/-3 y = 4/3x + 4